Chrome File Edit View History Bookmarks People Window Help xFIN3403.791518 tax rates are as show QUIZI- Caloulations x , Connect QUIZI Calculations Question 36 (of 38) 36, 10.00 points ooo points P6-3 Future Value and Multiple Cash Flows [LO1] Paradise, Inc., has identified an investment project with the following cash flows percent the future value of these cash flows in year 4 is$ Round your answer to 2 decimal places. (o.g. 32.16) If the dcount rate is 10 Do not inelude the dollar sign () $625 1,075 1,175 1,200 References Book &Resources Worksheet Difficulty: Basic P6-3 Future Value and Learning Objective: 06-1 12 名。 8 4 9
media%2Fc3f%2Fc3f5cb94-9567-4539-8cdc-b8
media%2Fba9%2Fba9296f4-d68f-4be4-9c01-43
Answer #1

Answer:

P6-3 Future value and multiple cash flows

Future value = present value (1+r)^t

Present value in year 4 =$1200

Rate of interest = 10%=0.1

Number of years = 4

Future value=$1200(1+0.1)^4

=$1200(1.1)^4=$1756.92

P6-12 calculation of EAR

EAR=(1+r)^t -1

Rate of interest 10% quarterly compounded

Rate of interest =2.5% annually

Number of periods = 4

EAR=(1+0.025)^4-1=10.38%

Rate of interest 16% compounding monthly

Rate of interest =1.33% annually

Number of periods =12

EAR=(1+0.013)^12-1

=(1.013)^12-1=1.1676-1=16.76%

Rate of interest rate 18% compounding daily

Interest rate=0.049%

Number of periods =365

EAR=(1+0.00049)^365-1=(1.00049)^365-1

=1.1958-1=19.58%

Interest rate 10% compounding infinitely

EAR=e^r-1

=e^0.1-1=2.71828^0.1-1

=0.105171=10.51%

P6-28 discounted cash flow analysis

Interest rate = 10%

Year 1

cash flow =$1500

Present value factor = 0.9090[1÷(1+r)^n]

Discounted cash flow =cash flow × PV factor

=$1500×0.9090=$1363.63

Year 2

Cash flows =0

Discounted cash flow =0

Year 3

Cash flows = $7200

Present value factor =0.7513[1÷(1+r)^n]

Discounted cash flow=cash flow×pv factor

=$7200×0.7513=$5,409

Year 4

Cash flows=900

Present value factor=0.6830[1÷(1+r)^n]

Discounted cash flow=cash flow×pv factor

=$900×0.6830=$614.71

Present value of cashflow's = $1363.63+0+$5409+$614.71=$7387.34